# Enhancing reliability in prediction intervals using point forecasters: Heteroscedastic Quantile Regression and Width-Adaptive Conformal Inference

### v1

@@ Abstract
Building prediction intervals for time series forecasting problems presents a complex challenge, particularly when relying solely on point predictors, a common scenario for practitioners in the industry. While research has primarily focused on achieving increasingly efficient valid intervals, we argue that, when evaluating a set of intervals, traditional measures alone are insufficient. There are additional crucial characteristics: the intervals must vary in length, with this variation directly linked to the difficulty of the prediction, and the coverage of the interval must remain independent of the difficulty of the prediction for practical utility. We propose the Heteroscedastic Quantile Regression (HQR) model and the Width-Adaptive Conformal Inference (WACI) method, providing theoretical coverage guarantees, to overcome those issues, respectively. The methodologies are evaluated in the context of Electricity Price Forecasting and Wind Power Forecasting, representing complex scenarios in time series forecasting. The results demonstrate that HQR and WACI not only improve or achieve typical measures of validity and efficiency but also successfully fulfil the commonly ignored mentioned characteristics.
@@ 1 Context of the problem
Machine learning’s application in critical decision-making has expanded dramatically, reaching into domains as vital as medical diagnostics (Kourou et al., 2015; Papadopoulos, 2011) and autonomous vehicle navigation (Melotti et al., 2023). The imperative to accurately assess the risks associated with model predictions in these areas cannot be overstated. This necessity is similarly crucial in managing time series data, which plays a significant role in various industries. For example, uncertainty quantification helps optimize inventory management by minimizing costs and preventing stock shortages (Böse et al., 2017). Additionally, in the energy sector, the integration of fluctuating renewable energy sources into the power grid not only complicates grid maintenance but also amplifies market volatility (Zhang et al., 2014; Nowotarski and Weron, 2018). Thus, the precise quantification of uncertainty becomes essential for informed decision-making using predictive models.
Additionally, it is typical, particularly in industrial settings, to not only develop in-house prediction models, if they are developed at all, but also to hire various forecasting services to provide predictions for the same process of interest. This approach allows companies to cross-validate the accuracy and reliability of different forecasts, thereby enhancing decision-making processes and optimizing operational efficiency. By comparing multiple forecasts, industries can mitigate risks associated with reliance on a single predictive model and gain a more comprehensive understanding of potential future scenarios.
Let us denote by $y_{t}$ the value of the variable whose evolution we wish to predict, which has been observed in the past at equidistant time points, $t=1,2,\ldots,T$. Suppose we have had access to predictions made by $M$ experts who independently provide point forecasts for each time point $t$, namely $\hat{y}_{t,1},\hat{y}_{t,2},\dots,\hat{y}_{t,M}$. We are interested in predicting the value of the variable at a future time point, $T+1$, that is $y_{T+1}$, based on the forecasts $\hat{y}_{T+1,1},\hat{y}_{T+1,2},\dots,\hat{y}_{T+1,M}$, and in providing a prediction interval $C_{\alpha,T+1}=\left[l_{\alpha,T+1},u_{\alpha,T+1}\right]$ such that
@@
$\mathbb{P}\left(y_{T+1}\in C_{\alpha,T+1}\right)=1-\alpha$
where $\alpha\in\left(0,1\right)$ represents the desired level of miscoverage, i.e., the proportion of times the band is allowed to not contain the actual value.
Let the length of the interval be $\left|C_{\alpha,T+1}\right|=u_{\alpha,T+1}-l_{\alpha,T+1}$. If for the same level of miscoverage $\alpha$ and two different instants $T_{1}$ and $T_{2}$, $\left|C_{\alpha,T_{1}+1}\right|>\left|C_{\alpha,T_{2}+1}\right|$ then one has a higher uncertainty at the instant $T_{1}$ than at $T_{2}$. This way, decisions taken at different points in time will be supported by a quantification of the risk that allows for a better understanding of the situation.
An interval $C_{\alpha,T+1}$ is considered valid if its marginal coverage is greater or equal than the objective coverage level $1-\alpha$. That is
$\mathbb{P}\left(y_{T+1}\in C_{\alpha,T+1}\right)\geq 1-\alpha.$
The efficiency of the interval is associated with the length of the interval. At the same level of coverage, the smaller length of the interval will be preferred, as it will better describe the process.
These two properties are the main ones when building an interval: the most efficient possible valid interval is desired (Shafer and Vovk, 2008). In practice, it is not feasible to assess the validity and efficiency of a prediction interval for a single time point. Typically, the analysis involves a comprehensive review of the intervals constructed over an extended period, which encompasses multiple prediction intervals. This broader evaluation helps to determine the overall effectiveness and reliability of the prediction intervals across different scenarios and conditions. Such an approach ensures a more solid understanding of how well the intervals perform in capturing the true values of the predicted variable under different circumstances. A good solution will give intervals of varying amplitude, where this variation should be related to the difficulty of the instance to be predicted (Angelopoulos and Bates, 2021). Thus, the variation of the length of the intervals must be measured and, in addition, it must be checked that there is an increasing relationship between the length of the interval and the error made by the point forecaster, because it would be expected to have a lower prediction accuracy in the most complicated situations. In this way, the intervals will be said to be informative.
Additionally, the validity of prediction intervals should not depend on the ease of prediction and, as previously argued, the complexity of making a prediction should be directly correlated with the length of the interval. For a given prediction interval $C_{\alpha,T+1}=\left[l_{\alpha,T+1},u_{\alpha,T+1}\right]$ with length $\left|C_{\alpha,T+1}\right|$, the objective is to transform this interval into a new one, $C^{*}_{\alpha,T+1}=\left[l^{*}_{\alpha,T+1},u^{*}_{\alpha,T+1}\right]$, such that the coverage probability remains independent of the length $\left|C_{\alpha,T+1}\right|$. This is mathematically represented as:
$\mathbb{P}\left(y_{T+1}\in C^{*}_{\alpha,T+1}|\left|C_{\alpha,T+1}\right|\right)=\mathbb{P}\left(y_{T+1}\in C_{\alpha,T+1}\right).$
This formulation ensures that confidence in the predictions remains constant, regardless of the length of the initial interval, thereby providing uniform certainty across situations with different difficulties.
These two concepts depend on the error of the point forecaster used. Thus, different point forecasters may give different error metrics. From now on, the point forecaster considered through out this work will be the mean of the $M$ different forecasts received, which is common practice and usually yields quite decent results despite its simplicity (Wang et al., 2023).
In this paper we focus on improving the results related to these properties that are key to make decisions supported by predictive models, while trying not to worsen the classical validity and efficiency measures in the context of probabilistic forecasting. All of this under the context in which only different (few) predictions of the process of interest are known, which is common in the industry.
@@
    We propose a quantile regression model, inspired by the philosophy of the Quantile Regression Aveaging (QRA) model of Nowotarski and Weron (2015), but with modifications so that there is an increasing relationship between the length of the interval and the difficulty of the prediction. Due to the particular use of the standard deviation of the point predictors, we will call the model Heteroscedastic Quantile Regression (HQR).
@@
    To provide theoretical coverage guarantees and to achieve uniformity of coverage regardless of the difficulty of the prediction, the Width-Adaptive Conformal Inference (WACI) method is proposed, which modifies the Adaptive Conformal Inference (ACI, Gibbs and Candès (2021)) method by solving the problems that the rest of the models may present in this regard.
To the best of our knowledge these two desired properties have not been considered before.
@@
The rest of the paper is structured in the following way. Section 2 discusses the different works related to uncertainty quantification in various forms and goes into detail on some of them, as they are the basis of the different contributions of the paper. Section 3 details our proposal, which is evaluated with the examples presented in Section 4 and whose discussion is extended in Section 5. Conclusions and future work finish the paper in Section 6.
2 Prior work
@@
Bayesian methods, by their very nature, are clear candidates for probabilistic regression problems. Through Bayes’ theorem, a posterior distribution can be obtained by updating beliefs as new information is obtained. Assuming a parametric model dependent on weights on the target variable, a distribution over these weights can be adopted. This is the approach followed in Bayesian neural networks (Neal, 2012). One can also consider the Bayesian approach directly on the target variable in the variant known as evidential regression (Amini et al., 2020) or with a functional approach through Gaussian processes (Rasmussen, 2003). However, Bayesian methods present problems that cannot be ignored, such as the choice of the prior distribution or the computational complexity.
@@
Assuming a specific distribution, one can try to estimate the distribution of $y_{T+1}$ based on the information known at time $T$. This is done by methods such as NGBoost (Duan et al., 2020), GAMLSS (Stasinopoulos and Rigby, 2008) as well as distributional neural networks and mixture density networks (Bishop, 1994). But as discussed in the previous section, the constraint of selecting a particular distribution can be quite restrictive.
@@
From a non-parametric point of view, classical methods such as bootstrapping the residuals to generate prediction intervals can be applied (Efron, 1987). However, the generality of the method tends not to produce the most satisfactory results. The application of quantile regression (Koenker and Bassett Jr, 1978) is also very popular, either through a linear model or by extending the method to more complex approaches such as neural networks (Cannon, 2011).
@@
All these methods can be easily extended to time series problems (for example by considering autoregressive effects, which is common practice) but none of them can assure the marginal coverage needed to provide valid prediction intervals. The Conformal Prediction framework (Vovk et al., 2005) ensures such marginal coverage in finite samples by assuming exchangeability between observations and without any assumptions about the probability distribution. In fact, Conformalized Quantile Regression (CQR) (Romano et al., 2019) extends quantile regression by providing the property of validity under exchangeability. However, as the exchangeability property is very demanding in time series, a large branch of research has focused on maintaining the good properties of the conformal predictors without assuming it. See for example (Gibbs and Candès, 2021; Zaffran et al., 2022; Gibbs and Candès, 2022; Bhatnagar et al., 2023).
@@
Regarding the context of the problem at hand, where only different predictors of the event to be forecasted are known, all these methodologies could be perfectly adapted by taking these predictors as explanatory variables. However, to our knowledge, there is only one work that has approached it in such a way: the Quantile Regression Averaging (QRA) model proposed by Nowotarski and Weron (2015), which will be explored in more depth in Section 2.1.1.
@@ 2.1 Quantile regression
As the initial objective should always be to obtain valid intervals, the estimation of quantiles, by its very definition, is a logical approximation to the problem. Let $\alpha\in(0,1)$ be the target miscoverage value, $\xi\in(0,\alpha)$ and let’s denote by $q_{\beta}(y_{T+1})$ the quantile of level $\beta$ of the distribution function $\mathcal{F}_{T+1}$ If we take $l_{\alpha,T+1}=q_{\xi}(y_{T+1})$ and $u_{\alpha,T+1}=q_{1-\alpha+\xi}(y_{T+1})$, then
$\mathbb{P}\left(y_{T+1}\in C_{\alpha,T+1}\right)\geq 1-\alpha.$
The typical choice for $\xi$ is $\xi=\frac{\alpha}{2}$ and this is the approach that will be followed in this paper. However, it should be noted that if the smallest intervals are desired, this option is not necessarily optimal.
Let $\bm{y}=\left(y_{1},y_{2},\dots,y_{n}\right)$ be the vector of $n$ observations of the variable of interest and $\bm{x}_{i}=(x_{i,1},x_{i,2},\dots,x_{i,m})$ be the vector of $m$ explanatory features for the observation $i$ for all $i\in\left\{1,2,\dots,n\right\}$. Let’s consider the model [^1] Without loss of generality, we consider the linear model.
$q_{\beta}(y_{i})=\lambda_{0}(\beta)+\lambda_{1}(\beta)x_{i,1}+\lambda_{2}(\beta)x_{i,2}+\dots+\lambda_{m}(\beta)x_{i,m}+\varepsilon_{i}(\beta);\;\mathbb{E}\left[\varepsilon_{i}(\beta)\right]=0$
@@
Let $y_{i}\sim Y$ and $\bm{x}_{i}\sim X$ for all $i=1,2,\dots,n$. Let $Y|X$ be distributed as $\mathcal{F}$. Just as the mean squared error is the loss function that is optimised to obtain the parameters that estimate the conditional mean as a point estimator, the estimation of the conditional quantile is achieved by optimising the pinball loss:
$l_{\beta}(y_{i},\hat{y}_{i})=\beta|y_{i}-\hat{y}_{i}|\mathds{1}\left\{y_{i}-\hat{y}_{i}\geq 0\right\}+(1-\beta)|y_{i}-\hat{y}_{i}|\mathds{1}\left\{y_{i}-\hat{y}_{i}\leq 0\right\}$
@@
$\hat{\bm{\lambda}}=\min_{\bm{\lambda}}\left\{\sum_{i=1}^{n}l_{\beta}(y_{i},\bm{\lambda}\cdot\bm{x}_{i})\ \right\}$
@@
$\hat{q}_{\beta}(y_{n+1})=\hat{\lambda}_{0}\left(\beta\right)+\hat{\lambda}_{1}\left(\beta\right)x_{n+1,1}+\hat{\lambda}_{2}\left(\beta\right)x_{n+1,2}+\dots+\hat{\lambda}_{m}\left(\beta\right)x_{n+1,m}$

### v2

@@ Abstract
Constructing prediction intervals for time series forecasting is challenging, particularly when practitioners rely solely on point forecasts. While previous research has focused on creating increasingly efficient intervals, we argue that standard measures alone are inadequate. Beyond efficiency, prediction intervals must adapt their width based on the difficulty of the prediction while preserving coverage regardless of complexity. To address these issues, we propose combining Heteroscedastic Quantile Regression (HQR) with Width- Adaptive Conformal Inference (WACI). This integrated procedure guarantees theoretical coverage and enables interval widths to vary with predictive uncertainty. We assess its performance using both a synthetic example and a real world Electricity Price Forecasting scenario. Our results show that this combined approach meets or surpasses typical benchmarks for validity and efficiency, while also fulfilling important yet often overlooked practical requirements.
@@ 1 Context of the problem
Probabilistic forecasting for time series is a key tool in fields such as finance, energy, and operations management, where understanding the range of potential future outcomes is as important as predicting the most likely value for optimal decision-making (Gneiting and Katzfuss, 2014). Given a time series $y_{1},y_{2},\dots,y_{T}$, $y_{t}\in\mathbb{R}$, the objective is to forecast the next $h$ steps, $y_{T+1},y_{T+2},\dots,y_{T+h}$. Unlike point forecasts that provide a single estimated value, probabilistic forecasting aims to model the conditional density $\mathbb{P}\left(y_{T+i}\left|y_{1},y_{2},\dots,y_{T},X_{1},X_{2},\dots,X_{T+i}\right.\right)$, for each of the next $h$ steps, where $X_{t}=\left(x_{t,1},x_{t,2},\dots,x_{t,k}\right)$ represents a vector of $k$ regressors available at each time step $t$.
The importance of probabilistic forecasting is particularly evident in high-stakes applications. For example, this can be seen in the electricity market through electricity price forecasting or renewable energy prediction (Zhang et al., 2014; Nowotarski and Weron, 2018). In the first case, high price volatility requires the presence of risk-aware strategies. Underestimating future prices may result in missed opportunities for higher profits, while overestimating prices could lead to losses from overcommitment or uncompetitive bidding. Probabilistic forecasts allow market participants to incorporate these cost asymmetries into their strategies. In renewable energy forecasting, the inherent unpredictability of weather conditions creates a pressing need for probabilistic methods. The power output of renewable sources like wind turbines and solar panels depends heavily on stochastic variables such as wind speed or solar irradiance. Probabilistic forecasts enable grid operators to account for this variability by scheduling backup generation and maintaining reserve capacity. They also help renewable energy producers avoid penalties by providing a clearer picture of the likelihood of over or under-delivery in their market bids.
Let’s assume that only one-step predictions are made ($h=1$) and that the conditional distribution to be modelled is unimodal. In this work, uncertainty is represented through prediction intervals, which provide a range within which future values are expected to fall with a specified probability. That is, given a miscoverage rate $\alpha\in\left(0,1\right)$ a prediction interval $\widehat{C}_{\alpha}\left(X_{T+1}\right)=\left[\widehat{l}_{\alpha}\left(X_{T+1}\right),\widehat{u}_{\alpha}\left(X_{T+1}\right)\right]\subseteq\mathbb{R}$ is build such that
$\mathbb{P}\left(y_{T+1}\in\widehat{C}_{\alpha}\left(X_{T+1}\right)\right)\geq 1-\alpha.$ (1)
An interval is said to be valid when property (1) is satisfied, i.e. when its marginal coverage is greater than or equal to the target coverage determined by the user. However, coverage alone does not guarantee optimal intervals. When building a prediction interval, the most efficient valid interval possible is desired (Shafer and Vovk, 2008). The efficiency of an interval is related to its length. When two prediction intervals achieve the same coverage level, the one with shorter length is preferred. This preference stems from the fact that narrower intervals provide more precise and actionable information, reducing the uncertainty range while maintaining reliability. Achieving the specified level of coverage with minimal interval length is crucial for constructing intervals that are not only statistically valid but also practically useful. The interval length of a prediction interval $\widehat{C}_{\alpha}\left(X_{T+1}\right)$ is denoted by $\left|\widehat{C}_{\alpha}\left(X_{T+1}\right)\right|$.
While these are crucial properties, this work asserts that they alone are not sufficient for ensuring practical utility. In addition to these considerations, we propose two further essential properties that prediction intervals must satisfy to enhance their applicability in real-world decision-making.
  1. 1.
    Adaptivity and correlation with prediction difficulty: prediction intervals should adapt to different levels of uncertainty present in the data. Specifically, the length of the intervals should be correlated with the difficulty of the prediction, such that shorter intervals are associated with easier-to-predict situations and longer intervals with more challenging ones. This ensures that the intervals effectively reflect the underlying uncertainty in the forecasting process.
  2. 2.
    Independence between coverage and interval length: In practice, the assessment of a prediction interval’s validity and efficiency requires a comprehensive evaluation over a period involving multiple prediction intervals. The empirical coverage is defined as
    $\frac{1}{N}\sum_{t=T}^{T+N}\mathds{1}\left(y_{t+1}\in\widehat{C}_{\alpha}\left(X_{t+1}\right)\right)$
    where $\mathds{1}\left(\cdot\right)$ is the indicator function and $N\in\mathbb{N}$ is the number of predictions that have been made. The marginal coverage of $\widehat{C}_{\alpha}\left(X_{T+1}\right)$ is approximated by this quantity. That is,
    ^$\widehat{\mathbb{P}}\left(y_{t}\in\widehat{C}_{\alpha}\left(X_{t}\right)\right)=\frac{1}{N}\sum_{t=T}^{T+N}\mathds{1}\left(y_{t+1}\in\widehat{C}_{\alpha}\left(X_{t+1}\right)\right).$
    Let $\mathcal{I}_{\rho}$ be the set of indices such that the length of the interval associated with that index is within $\delta>0$ of $\rho\in\mathbb{R}$, i.e,
    $\mathcal{I}_{\rho}=\left\{t+1\,:\,\left|\widehat{C}_{\alpha}\left(X_{t+1}\right)-\rho\right|\leq\delta\,,\,\rho\in\mathbb{R}\right\}.$
    Let $N_{\rho}$ be the number of elements of that set. For all $\rho$ such that $N_{\rho}\neq 0$, the desired property that we advocate is
@@
    ^$\widehat{\mathbb{P}}\left(y_{t}\in\widehat{C}_{\alpha}\left(X_{t}\right)\left|\left|\widehat{C}_{\alpha}\left(X_{t}\right)\right|\right.\approx\rho\right)=\frac{1}{N_{\rho}}\sum_{i\,\in\,\mathcal{I}_{\rho}}\mathds{1}\left(y_{i}\in\widehat{C}_{\alpha}\left(X_{i}\right)\right)=1-\alpha$ (2)
    for a specified miscoverage value $\alpha\in\left(0,1\right)$. This ensures that the coverage level is consistent and independent of the interval’s length, maintaining its reliability across different prediction difficulties.
While the first of these properties has been recognised in numerous papers as a desirable characteristic (Angelopoulos et al., 2023), to our knowledge there is only one paper (Feldman et al., 2021) that has dealt with the second and with a different perspective. In Feldman et al. (2021), independence between the coverage indicator and the interval length is pursued to enhance the conditional coverage of a quantile regression process. This approach is grounded in the observation that, for the true quantiles of the conditional distribution, the coverage indicator and the interval length are orthogonal. In this work this approximation is extended to the case of time series, but with the similar objective of making a better approximation of the real quantiles, thus avoiding biases in situations of low or high uncertainty.
Moreover, these properties are intrinsically linked to aleatoric and epistemic uncertainty. Adaptability to more or less predictable situations is directly tied to aleatoric uncertainty, as it ensures that prediction intervals accurately reflect variations in inherent randomness. For instance, periods of higher or lower volatility can be effectively differentiated. In contrast, the independence between coverage and interval length is associated with epistemic uncertainty, as it indicates the absence of biases linked to prediction difficulty. A method that produces noticeably different coverage levels depending on the interval length exhibits signs of inadequate modelling or insufficient data to capture the true data-generating process.
Finally, this work addresses a practical scenario commonly encountered in the industry, where only $M$ point forecasting models are available, and no additional information about the underlying data-generating process or the models themselves is accessible. Specifically, at time $T$, the only available information consists of the $M$ predictions $\bm{\widehat{y}}_{T+1}=\left(\widehat{y}_{T+1,1},\,\widehat{y}_{T+1,2},\,\dots,\,\widehat{y}_{T+1,M}\right)$ for $y_{T+1}$ along with their historical values. This situation is typical in contexts where organizations rely on external forecasting tools without detailed knowledge of their construction or assumptions. The primary goal of this work is to propose methods for generating reliable prediction intervals using only the outputs of these point forecasters, i.e, we want to build a prediction interval $\widehat{C}_{\alpha}(X_{T+1})\equiv\widehat{C}_{\alpha}(\bm{\widehat{y}}_{T+1})$, at the same time that the resulting intervals exhibit the two key properties introduced earlier, which are often overlooked in the literature. By addressing this limitation, the proposed approach ensures both theoretical soundness and practical applicability in industrial settings, making it a versatile solution for real-world forecasting challenges. To simplify the notation, in the rest of the paper the constructed interval will be denoted by $\widehat{C}_{\alpha,T+1}$, although it should be noted that the source of information in the construction comes from the different predictions.
@@
    A quantile regression model is proposed, inspired by the philosophy of the Quantile Regression Averaging (QRA) model of Nowotarski and Weron (2015), but with modifications so that there is an increasing relationship between the length of the interval and the difficulty of the prediction. Due to the particular use of the standard deviation of the point predictors, the model is called Heteroscedastic Quantile Regression (HQR).
@@
    To provide theoretical coverage guarantees and to achieve uniformity of coverage regardless of the difficulty of the prediction, the Width-Adaptive Conformal Inference (WACI) method is proposed, which modifies the Adaptive Conformal Inference (ACI, Gibbs and Candès (2021)) method by solving the problems that the rest of the models in the literature may present in this regard.
The combination of HQR with WACI ensures strong results in terms of both validity and efficiency while simultaneously fulfilling the two desired properties.
@@
The rest of the paper is structured in the following way. Section 2 discusses the different works related to uncertainty quantification in various forms and goes into detail on some of them (Sections 2.2, and 2.3), as they are the basis of the different contributions of the paper. Section 3 details our proposal, distinguishing HQR (Section 3.1) and WACI (Section 3.2), which is evaluated with a synthetic example presented in Section 4.2 and with a real life electricity price forecasting example in Section 4.3. Conclusions and future work finish the paper in Section 5.
2 Prior work on probabilistic forecasting
2.1 General overview
@@
Bayesian methods, by their very nature, are clear candidates for probabilistic prediction. Through Bayes’ theorem, a posterior distribution can be obtained by updating beliefs as new information is obtained. Assuming a parametric model dependent on weights on the target variable, a distribution over these weights can be adopted. This is the approach followed in Bayesian neural networks (Neal, 2012). One can also consider the Bayesian approach directly on the target variable in the variant known as evidential regression (Amini et al., 2020) or with a functional approach through Gaussian processes (Rasmussen, 2003). However, Bayesian methods present problems such as the choice of the prior distribution or the computational complexity.
@@
Assuming a specific distribution, one can try to estimate the distribution of $y_{T+1}$ based on the information known at time $T$. This is done by methods such as NGBoost (Duan et al., 2020), GAMLSS (Stasinopoulos and Rigby, 2008) as well as distributional neural networks and mixture density networks (Bishop, 1994). But the constraint of selecting a particular distribution can be quite restrictive. Data behaviour often evolve over time, and a fixed distribution may fail to remain valid as these changes occur. Additionally, it is common to assume simple distributions (e.g., the normal distribution), which often fail to capture the complex characteristics exhibited by real-world data, such as skewness, heavy tails, or multimodality. Conversely, selecting an overly complex distribution can lead to issues such as overfitting or excessive computational demands.
@@
From a non-parametric point of view, classical methods such as bootstrapping the residuals to generate prediction intervals can be applied (Efron, 1987). However, the generality of the method tends not to produce the most satisfactory results. The application of quantile regression (Koenker and Bassett Jr, 1978) is also very popular, either through a linear model or by extending the method to more complex approaches such as neural networks (Cannon, 2011). Quantile regression is explained in more detail in Section 2.2
@@
All these methods can be easily extended to time series problems (for example by considering autoregressive effects, which is common practice) but none of them can assure the marginal coverage needed to provide valid prediction intervals. The Conformal Prediction framework (Vovk et al., 2005) ensures such marginal coverage in finite samples by assuming exchangeability between observations and without any assumptions about the probability distribution. In fact, Conformalized Quantile Regression (CQR) (Romano et al., 2019) extends quantile regression by providing the property of validity under exchangeability, while trying to fit the heteroscedasticity properties encountered in the data. Indeed, the properties that we want to be found in the intervals to be constructed are closely tied to achieving conditional coverage, a topic of significant interest in the field of conformal predictions (Romano et al., 2019; Sesia and Romano, 2021; Chernozhukov et al., 2021; Han et al., 2022; Izbicki et al., 2022). However, the existing body of work on conditional coverage does not fully align with the framework considered here, as it typically assumes a static data distribution over time, a condition that is not met in our context. As the exchangeability property is very demanding in time series, a large branch of research has focused on maintaining the good properties of the conformal predictors without assuming it. See for example (Gibbs and Candès, 2021; Zaffran et al., 2022; Bhatnagar et al., 2023; Auer et al., 2023; Gibbs and Candès, 2024). More details related with Conformal Prediction can be found in Section 2.3
@@
Regarding the context of the problem at hand, where only different predictors of the event to be forecasted are known, most methodologies can, in principle, be adapted by treating these predictors as explanatory variables. However, to our knowledge, there is only one work that has approached it in such a way: the Quantile Regression Averaging (QRA) model proposed by Nowotarski and Weron (2015). Additionally, there are methods designed to combine predictors of the mean to enhance point forecasts, which have been adapted for probabilistic forecasting in an online setting, such as the approach proposed by Gaillard et al. (2016) using the algorithm in Gaillard et al. (2014). This paper will focus on QRA, which will be examined in greater detail in Section 2.2.1.
@@ 2.2 Quantile regression
Let $F_{T+1}$ be the cumulative distribution function of the random variable $Y_{T+1}$. For a given probability level $\beta\in\left(0,1\right)$, the quantile $\beta$ of $Y_{T+1}$ is defined as
$q_{\beta}\left(Y_{T+1}\right)=\inf\left\{y\in\mathbb{R}\,:\,F_{T+1}\left(y\right)\geq\beta\right\}.$
When constructing prediction intervals, prioritizing their validity is crucial. Valid intervals are essential for making informed decisions and managing risk, as they guarantee the desired level of coverage. If the intervals do not align with the expected confidence level, their reliability is compromised, leading to a loss of trust in the predictions and diminishing their value as decision-making tools. Therefore, the estimation of quantiles is a logical approximation to the problem. Let $\alpha\in(0,1)$ be the target miscoverage value and $\xi\in(0,\alpha)$. If $C_{\alpha,T+1}=\left[l_{\alpha,T+1},u_{\alpha,T+1}\right]=\left[q_{\xi}(Y_{T+1}),q_{1-\alpha+\xi}(Y_{T+1})\right]$ is considered, then
$\mathbb{P}\left(y_{T+1}\in C_{\alpha,T+1}\right)\geq 1-\alpha,$
where $y_{T+1}$ is the actual observed value at time $T+1$.
The typical choice for $\xi$ is $\xi=\frac{\alpha}{2}$ and this is the approach that will be followed in this paper. However, it should be noted that if the smallest intervals are desired, this option is not necessarily optimal. For example, when the distribution of $Y_{T+1}$ is asymmetric, other choices of $\xi$ can provide shorter intervals without sacrificing coverage.
Let $\bm{y}=\left(y_{1},y_{2},\dots,y_{n}\right)$ be the vector of $n$ observations of the variable of interest and $\bm{x}_{i}=(x_{i,1},x_{i,2},\dots,x_{i,m})$ be the vector of $m$ explanatory features for the observation $i$ for all $i\in\left\{1,2,\dots,n\right\}$. Let $y_{i}\sim Y$ and $\bm{x}_{i}\sim X$ for all $i=1,2,\dots,n$. Let $Y|X$ be distributed as $F$. Let’s consider the model
$q_{\beta}(Y|\bm{x}_{i})=\lambda_{0}(\beta)+\lambda_{1}(\beta)x_{i,1}+\lambda_{2}(\beta)x_{i,2}+\dots+\lambda_{m}(\beta)x_{i,m}+\varepsilon_{i}(\beta);\;\mathbb{E}\left[\varepsilon_{i}(\beta)\right]=0$
@@
Just as the mean squared error serves as the loss function optimized to estimate the conditional mean as a point estimator, the conditional quantile is estimated by minimizing the pinball loss function (Koenker and Bassett Jr, 1978):
$\ell_{\beta}(y_{i},\widehat{y}_{i})=\beta|y_{i}-\widehat{y}_{i}|\mathds{1}\left\{y_{i}-\widehat{y}_{i}\geq 0\right\}+(1-\beta)|y_{i}-\widehat{y}_{i}|\mathds{1}\left\{y_{i}-\widehat{y}_{i}\leq 0\right\}$
@@
$\widehat{\bm{\lambda}}=\min_{\bm{\lambda}}\left\{\sum_{i=1}^{n}\ell_{\beta}(y_{i},\bm{\lambda}^{\text{T}}\bm{x}_{i})\ \right\}$
@@
$\widehat{q}_{\beta}(Y|\bm{x}_{n+1})=\widehat{\lambda}_{0}\left(\beta\right)+\widehat{\lambda}_{1}\left(\beta\right)x_{n+1,1}+\widehat{\lambda}_{2}\left(\beta\right)x_{n+1,2}+\dots+\widehat{\lambda}_{m}\left(\beta\right)x_{n+1,m}$